# MORPH: Shape-agnostic PDE Foundation Models

### v2

@@ 5 Conclusion
We introduced MORPH, a shape-agnostic foundation model for PDEs that unifies heterogeneous spatiotemporal data across 1D–3D domains, scalar/vector fields, and mixed resolutions within a single convolutional vision transformer backbone. Our Unified Physics Tensor Format (UPTF-7) standardizes layout across PDE datasets while preserving physical semantics of the observations. By combining component-wise convolutions, inter-field cross-attention, and 4D axial attention, MORPH learns locality, cross-physics dependencies, and long-range spatiotemporal structure in a compute-aware way, and scales across multiple model variants. These design choices, together with parameter-efficient LoRA adapters for finetuning, enable effective transfer from diverse pretraining corpora to downstream tasks. We evaluated MORPH against its standalone version, task-specific baselines and recent PDE foundation models. Beyond prediction accuracy, MORPH highlights pathways toward flexible and scalable PDE foundation models for scientific machine learning.
@@
We are committed to ensuring the reproducibility of our results. All pretraining, finetuning, evaluation, visualization, and normalization code, together with pretrained model checkpoints, are publicly released on GitHub and Hugging Face. The datasets used in this work are sourced from PDEgym, PDEBench, and The Well.
@@ A.1.4 Compressible Navier-Stokes (CNS) system
Numerical scheme: For advection/shock terms: HLLC + MUSCL (robust, shock-capturing, 2nd order). For viscous/diffusion terms: central differencing (smooth, consistent with Navier–Stokes viscosity). More details on the solver Toro et al. (1994); Van Leer (1997).
@@
BCs & ICs: Non-periodic Dirichlet (no-slip) boundaries BCs. Initial conditions $v_{0}$ and inhomogeneous forcing parameters $u$ are independently sampled from isotropic Gaussian random fields with truncated power-law power spectra.
@@
Numerical scheme: Finite-difference simulations implemented in PhiFlow Holl et al. (2020).
@@ A.2 The Well
Description: Two-species reaction–diffusion model whose scalar concentrations vary in space and time. It exhibits rich pattern formations and is used as a canonical test for nonlinear spatiotemporal dynamics and biological morphogenesis (Gray & Scott, 1984).
@@
Description: A multiphysics dataset coupling fluid dynamics and electromagnetism. MHD is a strongly nonlinear system in which coupled fluid–magnetic dynamics generate broadband turbulence, shocks and compressions, intermittent current sheets, and reconnection, yielding larger dynamics range of spatial and temporal scales. MHD modeling finds applications in understanding solar winds, interstellar medium, magnetospheres, fusion energy devices like tokamaks (Burkhart et al., 2020).
@@
Numerical scheme: Third-order hybrid ENO shock-capturing scheme (Cho & Lazarian, 2003), isothermal EOS and forced turbulence. The high-res data were generated at $256^{3}$ and anti-aliased downsampled to $64^{3}$ for the MHD_64 split.
@@
Numerical scheme: Density-Independent SPH (DISPH) in ASURA-FDPS; radiative processes via tabulated cooling/heating (Hirashima et al., 2023).
Data specifications: N=20, [^0] N=20{,}000 trajectories, $T=21$ time steps, $D\times H\times W=12\times 64\times 64$ spatial points, one vector field variables $(v)$ and three scalar field $(\rho,P,T)$. Out of 4 fields, we only take 3 independent fields $(v,P,T)$. We use single dataset file corresponding to $T_{0}=10,\rho_{0}=0.445,Z=0.1$.
@@ A.3.1 FNS-KF
BCs & ICs: Periodic boundary conditions. Initial velocity is set exactly as in NS-PwC (Herde et al., 2024) by prescribing a piecewise-constant vorticity $\omega_{0}$ on a $p\times p$ uniform partition (with $p=10$) and recovering a divergence-free velocity:
@@
Data specifications: N=20, [^0] N=20{,}000 trajectories, $T=21$ time steps, $H\times W=128\times 128$, one vector field variables $(v)$. We assemble all the ‘.nc’ files in the data directory into a single combined file Herde et al. (2024).
@@ B.1 Model details
B.2 Data Processing
B.2.1 Unified Physics Tensor Format (UPTF-7)
[…]
B.2.2 Custom data sharding logic.
Input: Ordered files FILES, autoregressive window ar_order, chunk size, world size $W$, rank, num workers $K$, worker_id, base_seed, epoch
Output: Stream of $(X,y)$ pairs for this sub-worker
Count samples.
Balance across sub-workers.
Seed RNG $\leftarrow$ base_seed ++ epoch;
Stream and assign.;
foreach $f\in\texttt{FILES}$ do
   Open $f$ and iterate in chunks of size chunk_size;
   For each chunk: shuffle with the epoch seed and prepare AR pairs $(X,y)$;
 foreach $(x_{i},y_{i})\in(X,y)$ do
    if global_idx $\geq T^{\ast}$ then
       stop
    if (global_idx $\bmod G$) == my_id then
       yield $(x_{i},y_{i})$; if this sub-worker has yielded $m$ samples then
          stop
    global_idx $\leftarrow\texttt{global\_idx}+1$;
B.3 Pretraining settings
  • Normalization: We adopt Reversible Instance Normalization (RevIN(Kim et al., 2021). For each dataset used in pretraining and fine-tuning, we compute and cache dataset-level normalization statistics (e.g., mean and standard deviation), and at evaluation time we invert the transformation by denormalizing model outputs with the same constants. This strategy eliminates the need to normalize every batch on-the-fly, thereby reducing computational overhead during pretraining, finetuning and inference.
@@ B.3 Pretraining settings
  • Balanced task sampling: For multi-dataset training, each dataset $i$ is assigned a sampling weight inversely proportional to its number of trajectories $N_{i}$, i.e., $w_{i}\propto 1/N_{i}$, and the weights are normalized so that $\sum_{i}w_{i}=1$.
@@
  • Parallelization: Single-GPU, multi-GPU on a single node via DataParallel, and multi-node, multi-GPU via DistributedDataParallel.
@@
  • Optimizer: We use AdamW (Loshchilov & Hutter, 2017) with a learning rate of $10^{-3}$ and weight decay $10^{-2}$. For standalone CFD datasets, we use $10^{-4}$. The learning-rate schedule is ReduceLROnPlateau with a decay factor of $0.5$ and a patience of $5$ epochs, applied after a $20$-epoch warmup.
@@
B.4 Finetuning settings
B.4.1 MORPH Models
@@
B.4.2 Existing SOTA Models
To contextualize MORPH against recent PDE foundation models, we fine-tune two Poseidon variants (Poseidon-T (21M) and Poseidon-B (158M) Herde et al. (2024)) and two DPOT variants (DPOT-S (30M) and DPOT-M (122M) Hao et al. (2024)). Results appear in the lowest block of Table 2 and in Table 4.
@@ B.4.2 Existing SOTA Models
For the rollout experiment (Table 4), we fine-tune on the first 128 trajectories for 200 epochs and evaluate on the last 240 trajectories of the FNS-KF subset of PDEBench, reporting average MSE and RMSE over 20-step rollouts Takamoto et al. (2022). Throughout, all inputs are at $128{\times}128$ spatial resolution per channel, and we apply the data normalization procedures specified by each repository.
@@
DPOT is pretrained with a 10-step temporal context at $128{\times}128$ resolution. To adapt DPOT for single-step $t\!\rightarrow\!t{+}1$ prediction (Table 2), we follow the input-padding strategy analogous to Herde et al. (2024), feeding 10 copies of the state at time $t$ to satisfy the context window:
@@
Poseidon takes a single input frame and a selectable $\Delta t$ and predicts $t{+}\Delta t$. For the single-step experiments (Table 2), we fine-tune Poseidon with $\Delta t{=}1$ on all $t\!\rightarrow\!t{+}1$ pairs in the training set. For full-trajectory prediction (Table 4), we adopt the repository’s default $\Delta t$ fine-tuning schedule with max_num_time_steps =7=7 and time_step_size =2=2, which enables multi-step forecasting. Results shown for Poseidon in Table 4 are that of Poseidon’s direct (non-autoregressive) prediction, where the model directly predicts every time-step in the trajectory without feeding any predictions back into the model. 2D-CFD data in Table 2 is of ($512\times 512$) resolution. To align with Poseidon’s pretrained weights, we downsample these images to $128\times 128$ with area averaging, run inference/training at $128\times 128$, then upsample back to ($512\times 512$) with bilinear interpolation for evaluation against ground truth. All Poseidon models were fine-tuned with AdamW (LR $1\times 10^{-4}$), batch size 40, cosine LR schedule, and gradient clipping (max grad norm =5.0=5.0).
@@ Appendix C Additional and Extended Results
C.1 Data and compute-scarce finetuning
The data and compute-scarce fine-tuning results for MORPH-FM-Ti on the 1D-DR dataset as shown in Fig. 2. The Fig. 2 reports RMSE as a function of the percentage of fine-tuning trajectories for the 1D-DR prediction task. We compare against the standalone MORPH-SS-Ti model trained on the full set of trajectories. Notably, the pretrained model outperforms the standalone model even when fine-tuned with only 25% of the trajectories.
Figure 2: MORPH-FM-Ti fine-tuning on 1D-DR: RMSE vs $\%$ different trajectories of 1D-DR dataset for 100 epochs.
In Fig. 3, we conduct an additional data- and compute-scarce study by fine-tuning MORPH-FM-S on the 2D-FNS-KF dataset. We report RMSE as a function of the number of fine-tuning trajectories. With just $\sim$ 256 trajectories (less than 1% of the full set) and only 100 fine-tuning epochs, the fine-tuned MORPH-FM-S already surpasses the standalone MORPH-SS-S trained on all trajectories.
Figure 3: MORPH-FM-S fine-tuning on 2D-FNS-KF: RMSE vs # trajectories of 2D-FNS-KF dataset for 100 epochs.
[…]
@@ C.1 Data and compute-scarce finetuning
C.2 Autoregressive rollouts
Figs. 4 and 5 show 10-step autoregressive rollouts for MORPH-SS-Ti and MORPH-SS-S on the Shallow Water Equations (SWE) dataset, using the t=0t{=}0 snapshot as input. Across the rollout horizon, MORPH-SS-S exhibits lower error than MORPH-SS-Ti, consistent with the NRMSE reported in Table 2. Both models produce stable multi-step forecasts, exhibiting limited error accumulation and no blow-up for the 10-step horizon.

### v3

@@ 4.3 Scaling studies
4.4 Ablation Studies
Model architecture.
Table 6 summarizes ablations over the convolutional operator, field-fusion module, and attention operator in MORPH. Introducing a lightweight convolutional operator with 8 filters consistently reduces error on 2D/3D benchmarks with only a modest increase in GFLOPs, while having negligible effect on 1D. For field fusion, cross-attention outperforms a simple concat+MLP scheme across datasets, at nearly identical parameter counts and modest additional compute. Among attention variants, axial attention systematically offers the best accuracy–compute trade-off, outperforming both full and sparse self-attention. The gains are most pronounced at larger patch-token counts, where axial factorization yields large FLOP savings while still improving test loss. More details are deferred to Appendix C.2.
Table 6: Ablation studies (Model architecture): Effect of convolution operator, cross-attention field fusion and axial attention on the total number of parameters (in M), floating point operations of the operator (in GFLOPs), and test loss (MSE) for different datasets. ps = patch size, Np - number of patches or tokens.
Ablations Variant MHD3D DR2D CFD1D
Param. GFLOPs Loss Param. GFLOPs Loss Param. GFLOPs Loss
Conv Filters 0 (no conv) 8.29 0 0.18376 7.04 0 0.00787 7.14 0 0.17185
8 8.94 2.7556 0.16851 7.96 0.11376 0.00747 8.06 0.011 0.17348
Field-fusion Concat & Dense 8.8 0.2013 0.17101 7.89 0.10066 0.00827 7.99 0.05 0.18141
Cross-attention 8.94 0.4713 0.16851 7.96 0.16829 0.00747 8.06 0.1178 0.17348
Attention CFD2D-IC (ps = 8, Np = 4096) CFD2D-IC (ps = 16, Np = 1024) SW2D (ps = 8, Np = 256)
Full 5.78 5.9056 0.01228 14.2 75.141 0.01208 4.67 0.6711 0.00291
Sparse 5.78 1.8832 0.01324 14.2 8.6064 0.00856 4.67 0.4372 0.00208
Axial 8.94 1.8832 0.00812 17.3 8.6064 0.00724 7.83 0.4372 0.00172
Pretraining corpus.
To evaluate the effect of the pretraining set and to disentangle the influence of the dataset from that of the model architecture, we pretrain MORPH-FM-S*(27M) on the same datasets as Poseidon. We then finetune both MORPH and Poseidon on 128 trajectories and test on 240 trajectories for 200 epochs across three downstream tasks: CE-RM, NS-PwC, and NS-SL. The results reported in Table 9 (Appendix C.2) demonstrate that MORPH-FM-S*(27M), with only one-fifth the number of parameters, performs competitively with or even outperforms Poseidon. These findings indicate that the MORPH architecture yields performance gains that are largely independent of the choice of pretraining corpus. More details about this study are provided in Appendix C.2, and full autoregressive rollouts are shown in Figs. 10, 12 and 11 (Appendix C.3).
@@ 5 Conclusion
We introduced MORPH, a modality-agnostic foundation model for PDEs that unifies heterogeneous spatiotemporal data across 1D–3D domains, scalar/vector fields, and mixed resolutions within a single convolutional vision transformer backbone. Our Unified Physics Tensor Format (UPTF-7) standardizes layout across PDE datasets while preserving physical semantics of the observations. By combining component-wise convolutions, inter-field cross-attention, and 4D axial attention, MORPH learns locality, cross-physics dependencies, and long-range spatiotemporal structure in a compute-aware way, and scales across multiple model variants. These design choices, together with parameter-efficient LoRA adapters for finetuning, enable effective transfer from diverse pretraining corpora to downstream tasks. We evaluated MORPH against its standalone version, task-specific baselines and recent PDE foundation models. Beyond prediction accuracy, MORPH highlights pathways toward flexible and scalable PDE foundation models for scientific machine learning.
@@
All datasets used for pretraining and fine-tuning are publicly available. We will release the full codebase and model checkpoints under an open-source license upon publication.
@@ A.1.4 Compressible Navier-Stokes (CNS) system
Numerical scheme: For advection/shock terms: HLLC + MUSCL (robust, shock-capturing, 2nd order). For viscous/diffusion terms: central differencing (smooth, consistent with Navier–Stokes viscosity). More details on the solver toro1994restoration; van1997towards.
@@
BCs & ICs: Non-periodic Dirichlet (no-slip) BCs. Initial conditions $v_{0}$ and inhomogeneous forcing parameters $u$ are independently sampled from isotropic Gaussian random fields with truncated power-law power spectra.
@@
Numerical scheme: Finite-difference simulations implemented in PhiFlow (holl2020phiflow).
@@ A.2 Benchmark: The Well
Description: Two-species reaction–diffusion model whose scalar concentrations vary in space and time. It exhibits rich pattern formations and is used as a canonical test for nonlinear spatiotemporal dynamics and biological morphogenesis (gray1984autocatalytic).
@@
Description: A multiphysics dataset coupling fluid dynamics and electromagnetism. MHD is a strongly nonlinear system in which coupled fluid–magnetic dynamics generate broadband turbulence, shocks and compressions, intermittent current sheets, and reconnection, yielding larger dynamics range of spatial and temporal scales. MHD modeling finds applications in understanding solar winds, interstellar medium, magnetospheres, fusion energy devices like tokamaks (burkhart2020catalogue).
@@
Numerical scheme: Third-order hybrid ENO shock-capturing scheme (cho2003compressible), isothermal EOS and forced turbulence. The high-res data were generated at $256^{3}$ and anti-aliased downsampled to $64^{3}$ for the MHD_64 split.
@@
Numerical scheme: Density-Independent SPH (DISPH) in ASURA-FDPS; radiative processes via tabulated cooling/heating (hirashima20233d).
Data specifications: N=20, [^0] N=20{,}000 trajectories, $T=21$ time steps, $D\times H\times W=64\times 64\times 64$ spatial points, one vector field variable $(v)$ and three scalar field $(\rho,P,T)$. Out of 4 fields, we only take 3 independent fields $(v,P,T)$. We use single dataset file corresponding to $T_{0}=10,\rho_{0}=0.445,Z=0.1$.
@@ A.3.1 FNS-KF
BCs & ICs: Periodic boundary conditions. Initial velocity is set exactly as in NS-PwC (herde2024poseidon) by prescribing a piecewise-constant vorticity $\omega_{0}$ on a $p\times p$ uniform partition (with $p=10$) and recovering a divergence-free velocity:
@@
Data specifications: N=20, [^0] N=20{,}000 trajectories, $T=21$ time steps, $H\times W=128\times 128$, one vector field variables $(v)$. We assemble all the ‘.nc’ files in the data directory into a single combined file (herde2024poseidon). However, we use limited trajectories.
@@
A.4 Data Processing
Unified Physics Tensor Format (UPTF-7).
[…]
Normalization.
We use Reversible Instance Normalization (ReVIN) to counter covariate shift across heterogeneous datasets and training regimes. It was first proposed for time-series forecasting and validated to mitigate distribution shift (kim2021reversible). We normalize the data and cache the corresponding means and standard deviations, then use these statistics to exactly denormalize model outputs. Unlike mccabe2024multiple, which normalizes on-the-fly, we pre-normalize the entire dataset and train on normalized batches, reducing the normalization overhead incurred during training and fine-tuning, and inference.
Data streaming and sharding setup.
MORPH is designed to handle data heterogeneity and diversity, which introduces system-level challenges. To utilize CPU, GPU, and RAM efficiently, we make several non-traditional choices. One issue is that computing cluster resources impose RAM limits, making it difficult to load all datasets into memory simultaneously. As PDE benchmarks continue to grow, streaming becomes essential for building large PDE foundation models. Accordingly, we stream datasets from storage in chunks, which uses RAM efficiently and scales with the number and size of datasets. Another issue is the wide range of token lengths (128–4096) across pretraining datasets, which leads to varying computational burdens in parallel processing across batches. Instead of a single PyTorch DataLoader for all datasets, we use one dataloader per dataset. With six pretraining datasets, we run six DataLoaders. This enables dataset-specific choices of batch sizes and number of workers. Given this multi-dataloader setup, data sharding under DistributedDataParallel(DDP) is nontrivial, so we implement custom sharding across workers and ranks.
Input: Ordered files FILES, autoregressive window ar_order, chunk size, world size $W$, rank, num workers $K$, worker_i base_seed, epoch Output: Stream of $(X,y)$ pairs for this sub-worker
Count samples.
Balance across sub-workers.
Seed RNG $\leftarrow$ base_seed ++ epoch;
Stream and assign.;
foreach $f\in\texttt{FILES}$ do
   Open $f$ and iterate in chunks of size chunk_size;
   For each chunk: shuffle with the epoch seed and prepare AR pairs $(X,y)$;
 foreach $(x_{i},y_{i})\in(X,y)$ do
    if global_idx $\geq T^{\ast}$ then
       stop
    if (global_idx $\bmod G$) == my_id then
       yield $(x_{i},y_{i})$; if this sub-worker has yielded $m$ samples then
          stop
    global_idx $\leftarrow\texttt{global\_idx}+1$;
@@ B.1 Model details
B.2 Pretraining settings
Balanced task sampling.
We train MORPH on multiple pretraining datasets with varying numbers of trajectories per dataset. We perform balanced task sampling to avoid imbalanced learning and catastrophic forgetting. Each dataset is assigned a sampling weight inversely proportional to its number of trajectories, so datasets with fewer trajectories are randomly sampled more frequently. In our corpus, the 3D-MHD dataset has the fewest trajectories and is therefore sampled more often than the remaining five. For multi-dataset training, each dataset $i$ is assigned a sampling weight inversely proportional to its number of trajectories $N_{i}$, i.e., $w_{i}\propto 1/N_{i}$, and the weights are normalized so that $\sum_{i}w_{i}=1$. The resulting empirical per-batch sampling probabilities are approximately: 3D-MHD: 0.31, 2D-CFD-IC: 0.19, 3D-CFD: 0.18, 2D-DR: 0.12, 2D-SW: 0.12, and 1D-CFD: 0.08.
LoRA layers.
Low-Rank Adaptation (LoRA) is used for fine-tuning LLMs where the pretrained model weights are frozen while introducing trainable rank decomposition matrices into the transformer layers, dramatically reducing task-specific parameters during finetuning (hu2022lora). The motivation is supported by evidence that over-parameterized networks possess low intrinsic dimensionality, suggesting that task-specific updates lie in a low-rank subspace (aghajanyan2020intrinsic). Mathematically, for a frozen weight $W_{0}\in\mathbb{R}^{d\times k}$, LoRA parameterizes the update as
yielding the adapted weight
$W^{\prime}=W_{0}+\frac{\alpha}{r}\,BA,$
Compute infrastructure.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
Other details.
@@ B.2 Pretraining settings
  • Optimizer: We use AdamW (loshchilov2017decoupled) with a learning rate of $10^{-3}$ and weight decay $10^{-2}$. For standalone CFD datasets, we use $10^{-4}$. The learning-rate schedule is ReduceLROnPlateau with a decay factor of $0.5$ and a patience of $5$ epochs, applied after a $20$-epoch warmup.
@@
B.3 Finetuning settings
B.3.1 MORPH Models
@@
B.3.2 Existing SOTA Models
To contextualize MORPH against recent PDE foundation models, we fine-tune two Poseidon variants (Poseidon-T (21M) and Poseidon-B (158M) (herde2024poseidon)) and two DPOT variants (DPOT-S (30M) and DPOT-M (122M) (hao2024dpot)). Results are presented in Tables 5, 8 and 9.
@@ B.3.2 Existing SOTA Models
For the rollout experiment (Tables 8 and 9), we fine-tune on the first 128 trajectories for 200 epochs and evaluate on the last 240 trajectories of the FNS-KF subset of PDEBench, reporting average MSE and RMSE over 20-step rollouts (takamoto2022pdebench). Throughout, all inputs are at $128{\times}128$ spatial resolution per channel, and we apply the data normalization procedures specified by each repository.
@@
DPOT is pretrained with a 10-step temporal context at $128{\times}128$ resolution. To adapt DPOT for single-step $t\!\rightarrow\!t{+}1$ prediction (Table 5), we follow the input-padding strategy analogous to herde2024poseidon, feeding 10 copies of the state at time $t$ to satisfy the context window:
@@
Poseidon takes a single input frame and a selectable $\Delta t$ and predicts $t{+}\Delta t$. For the single-step experiments (Tables 5 and 9), we fine-tune Poseidon with $\Delta t{=}1$ on all $t\!\rightarrow\!t{+}1$ pairs in the training set. For full-trajectory prediction (Table 8), we adopt the repository’s default $\Delta t$ fine-tuning schedule with max_num_time_steps =7=7 and time_step_size =2=2, which enables multi-step forecasting. Results shown for Poseidon in Table 8 are that of Poseidon’s direct (non-autoregressive) prediction, where the model directly predicts every time-step in the trajectory without feeding any predictions back into the model. 2D-CFD data in Table 5 is of ($512\times 512$) resolution. To align with Poseidon’s pretrained weights, we downsample these images to $128\times 128$ with area averaging, run inference/training at $128\times 128$, then upsample back to ($512\times 512$) with bilinear interpolation for evaluation against ground truth. All Poseidon models were fine-tuned with AdamW (LR $1\times 10^{-4}$), batch size 40, cosine LR schedule, and gradient clipping (max grad norm =5.0=5.0).
@@ Appendix C Additional and Extended Results
C.1 Scaling studies
Scaling w.r.t. pretraining dataset size.
The scaling experiments are conducted on four H100 GPUs using Distributed Data Parallel (DDP). We use the per-GPU batch sizes (global batch size divided by 4) described in Sec. B.2. To vary the effective dataset size, we subsample the training data by randomly retaining a fraction $p\%$ of the mini-batches produced by the training dataloader and processing only these mini-batches in the model’s forward pass. Fig. 5 presents the study.
Figure 5: Scaling studies: Data-level scaling for MORPH-FM-S model
Scaling w.r.t. finetuning model size (across FMs).
We perform full fine-tuning of MORPH-FM-S using a learning rate of 1e-4, batch size 40, and weight decay 1e-4. We employ a ReduceLROnPlateau scheduler with a wait of 5 epochs and early stopping with a patience of 10 epochs. The hyperparameter settings and fine-tuning setup for Poseidon and DPOT are provided in Appendix B.3.2. Table 8 shows the comparison.
@@ C.1 Scaling studies
C.2 Ablation studies
Model architecture.
Table 6 evaluates three key design choices in MORPH i.e., the component-wise convolutional preprocessor, the field-fusion module, and the attention operator. We report parameter count (M), operator FLOPs (GFLOPs), and test MSE on representative datasets (MHD3D, DR2D, CFD1D, CFD2D-IC, and SW2D; here $ps$ denotes patch size and $N_{p}$ the number of patches/tokens). The ablation studies are performed on 4x H100s GPUs with same hyperparameters.
Convolutional preprocessor. Comparing “0 (no conv)” to the 8-filter variant shows that adding a lightweight convolutional operator consistently reduces error on 3D and 2D problems with modest compute overhead: on MHD3D the loss drops from $0.18376$ to $0.16851$ (about $8.3\%$) at the cost of $2.76$ GFLOPs, and on DR2D from $0.00787$ to $0.00747$ (about $5.1\%$) with only $0.11$ additional GFLOPs. For CFD1D the effect is small and slightly negative ($0.17185\rightarrow 0.17348$), indicating that the convolutional inductive bias is most beneficial for higher-dimensional settings.
Field-fusion. For field fusion, cross-attention is compared against a simple concat+dense baseline. Cross-attention consistently improves accuracy with negligible parameter overhead and a moderate increase in FLOPs: on MHD3D the loss decreases from $0.17101$ to $0.16851$ ($\approx 1.5\%$), on DR2D from $0.00827$ to $0.00747$ ($\approx 9.7\%$), and on CFD1D from $0.18141$ to $0.17348$ ($\approx 4.4\%$), while GFLOPs roughly double but remain small in absolute terms. This supports the use of content-aware attention for fusing multiple fields instead of static mixing.
Attention operator. The bottom block compares axial, full, and sparse self-attention for different token counts. Axial attention offers the best accuracy–compute trade-off and its advantage grows with sequence length. On CFD2D-IC with ps=16ps{=}16 and $N_{p}{=}1024$, axial attention achieves a loss of $0.00724$ using $8.61$ GFLOPs, whereas full attention reaches $0.01208$ at $75.14$ GFLOPs (about $8.7\times$ more compute and $\approx 40\%$ higher error). For ps=8ps{=}8, $N_{p}{=}4096$, axial is again superior ($0.00812$ vs. $0.01228$) while being $\approx 3.1\times$ cheaper in FLOPs, and a similar pattern holds on SW2D where axial is $\approx 1.5\times$ cheaper and $\approx 41\%$ lower in loss than full attention. Compared to sparse attention, axial has essentially the same FLOPs but reduces loss by $15$-$40\%$ across datasets. These trends are consistent with the theoretical complexity reduction from $O((tDHW)^{2})$ for full attention to O(tDHW(t+D+H+W))O(tDHW\,(t{+}D{+}H{+}W)) for axial factorization, and motivate axial attention as a near–Pareto-optimal choice for large spatiotemporal token grids.
Pretraining corpus.
We pretrained MORPH-FM-S (30M) from scratch on the same pretraining corpus as Poseidon i.e., NS-Sines, NS-Gaussians, CE-RP, CE-CRP, CE-KH, CE-Gauss herde2024poseidon for 100 epochs. We employ a ReduceLROnPlateau scheduler with a wait of 5 epochs and early stopping with a patience of 10 epochs. We then finetune MORPH on 128 trajectories and test on 240 trajectories for 200 epochs across three downstream tasks: CE-RM, NS-PwC, and NS-SL. We use a learning rate of 1e-4, batch size 40, and weight decay 1e-4, early stopping with patience of 25 during the finetuning. The hyperparameter settings for Poseidon are presented in Appendix B.3.2. Table 8 shows the comparison. Full autoregressive rollouts are shown in Figs. 10, 12 and 11.
Table 9: Ablation studies (Pretraining corpus): MORPH-FM-S*(27M) and Poseidon-FM-B are finetuned with 128 trajectories for 200 epochs: Rollout MSE on test set (lower is better) average over 240 trajectories.
Models FT datasets CE-RM NS-PwC NS-SL
Poseidon-FM-B (158M) 0.4181 0.0004 0.0163
MORPH-FM-S* (27M) 0.09044 0.00408 0.0159
C.3 Autoregressive rollouts
C.3.1 Standalone surrogate
Figs. 6 and 7 show 10-step autoregressive rollouts for MORPH-SS-Ti and MORPH-SS-S on the Shallow Water Equations (SWE) dataset, using the t=0t{=}0 snapshot as input. Across the rollout horizon, MORPH-SS-S exhibits lower error than MORPH-SS-Ti, consistent with the NRMSE reported in Table 3. Both models produce stable multi-step forecasts, exhibiting limited error accumulation and no blow-up for the 10-step horizon.